OpenStudy (anonymous):

How to solve 2x^2-12=-5x

pooja195 (pooja195):

Subtract -5x from both sides. \[\huge~\rm~2x^2−12−−5x=−5x−−5x\] \[\huge~\rm~ 2x^2+5x−12=0\] From here we need to factor by grouping  \[\huge~\rm~ax^2+bx+c=0\] \[\huge~\rm~ 2x^2+5x−12=0\] Using the stander-ed form equation we determine what a and c are in this case they are a=2 b=-12 When we multiply them together we get -24 B remains as 5  What numbers add up to 5 but multiply to make -24?